# Feature Selection Library (MATLAB Toolbox)

### 21 Feb 2024 v7

@@ Abstract
The Feature Selection Library (FSLib) signifies a notable progression in machine learning and data mining for MATLAB users, emphasizing the critical role of Feature Selection (FS) in enhancing model efficiency and effectiveness by pinpointing essential features for specific tasks. FSLib’s contributions are comprehensive, tackling various FS challenges. It offers a wide array of FS algorithms, including filter, embedded, and wrapper methods, allowing for optimal feature selection tailored to specific problem requirements. Filter methods prioritize intrinsic feature properties, embedded methods integrate selection within the training process, and wrapper methods evaluate features based on model performance, catering to diverse modeling approaches. FSLib also addresses the curse of dimensionality by facilitating the selection of relevant feature subsets, thereby reducing data dimensionality, lessening computational demands, and potentially improving model generalizability. Furthermore, by eliminating superfluous features, FSLib streamlines the learning process, enhancing model training efficiency and scalability. This targeted selection process not only accelerates model development but also bolsters model accuracy, precision, and recall by concentrating on crucial information. Additionally, FSLib enhances data interpretability, offering insights into data structure through the identification of significant features, thereby aiding in pattern discovery and understanding. In essence, FSLib extends beyond simple feature selection, providing a comprehensive framework that augments the entire machine learning and data mining workflow. By presenting an extensive selection of algorithms, mitigating dimensional challenges, expediting learning, improving model metrics, and fostering data insight, FSLib emerges as an instrumental resource in the evolution of machine learning research and practice.
@@ 1 Introduction
Feature selection (FS) is a critical preprocessing step in machine learning, aimed at enhancing the performance and efficiency of classification models by reducing data dimensionality and eliminating noise. This process is instrumental in refining the dataset to include only the most informative features, thereby improving the accuracy and computational efficiency of classifiers. The essence of FS lies in its ability to filter out redundant or irrelevant information, a task that becomes increasingly challenging as data complexity grows. Traditionally, feature identification has been manually conducted by experts within various learning domains. However, this manual selection process may not always identify the most relevant or non-redundant features, highlighting the necessity for automatic feature selection methods. These automated processes are adept at discarding irrelevant, redundant, and noisy data, which is crucial for enhancing learning and classification performance. The importance of feature selection is well-acknowledged across numerous fields, including machine learning, artificial intelligence, computer vision, and data mining, finding application in information retrieval, user re-identification, recommendation systems, and visual object tracking, among others Roffo et al. (2013a, b); Cristani et al. (2012); Roffo et al. (2014a, b); Roffo and Vinciarelli (2016); Hua and Jégou (2016); Kristan et al. (2016); Roffo and Melzi (2016a, b); Vinciarelli et al. (2018); Roffo and Melzi (2016a).
@@
Within the FS domain, methodologies are broadly classified into three primary categories: filter methods, wrapper methods, and embedded methods. Filter methods prioritize the intrinsic properties of the data, operating independently of any classifier to select features. Wrapper methods involve optimizing a predictor as part of the selection process, often yielding superior results compared to filter methods but at a higher computational cost. Embedded methods integrate the feature selection process with the model learning mechanism, making the two inseparable. This approach is unique in its interaction between feature selection and learning, exemplifying the importance of the function class under consideration in determining feature relevance.
@@
Wrapper methods employ classifiers to assess the utility of feature subsets, embedded methods integrate selection within the classifier learning process, and filter methods analyze data’s intrinsic properties irrespective of the classifier Guyon (2006). These approaches can execute two primary operations: ranking and subset selection. Ranking evaluates the importance of individual features without considering their interactions, whereas subset selection identifies the optimal feature combination for use. Although these operations can be sequential or exclusive, subset selection is inherently supervised, contrasting with the potentially unsupervised nature of ranking Guyon et al. (2002); Bradley and Mangasarian (1998); Grinblat et al. (2010); Zaffalon and Hutter (2002); Liu and Motoda (2008); Yu et al. (2012); Gu et al. (2012).
@@
Embedded methods uniquely intertwine feature selection with the learning algorithm, making them inseparable from the learning process and thereby tailoring feature selection to the specific model in use, such as Support Vector Machines Guyon (2006). This integration of selection and learning processes within embedded methods is particularly relevant in the context of modern attention and self-attention mechanisms in transformers. These mechanisms inherently perform a form of feature selection by weighting the importance of different parts of the data (e.g., tokens in natural language processing), thus aligning with the principles of embedded feature selection methods but on a data-driven, adaptive basis.
@@
Feature selection is recognized as an NP-hard problem Guyon (2006), necessitating the use of heuristic or suboptimal search strategies to manage the combinatorial complexity of selecting an optimal subset from a potentially vast feature space. Filters approach this challenge by initially ranking features individually before extracting a subset, with notable examples including Inf-FS Roffo et al. (2015a), EC-FS Roffo and Melzi (2016c), MutInf Zaffalon and Hutter (2002), and Relief-F Liu and Motoda (2008). Conversely, wrapper and embedded methods evaluate and select feature subsets in a more holistic manner, with FSV Bradley and Mangasarian (1998); Grinblat et al. (2010) and SVM-RFE Guyon et al. (2002) serving as key examples. Given the NP-hard nature of feature selection, with the combinatorial explosion of feature subsets to evaluate, efficient and heuristic search strategies are employed Guyon (2006). Filters prioritize individual features through ranking before extracting a subset, exemplified by methods such as Inf-FS and MutInf, while wrappers and embedded techniques evaluate and select subsets directly, with notable examples including FSV and SVM-RFE Roffo et al. (2015a); Roffo and Melzi (2016c); Zaffalon and Hutter (2002); Liu and Motoda (2008); Bradley and Mangasarian (1998); Grinblat et al. (2010); Guyon et al. (2002); Anderson et al. .
@@
Feature selection algorithms are also distinguished by their learning context, being either supervised or unsupervised. Supervised algorithms leverage class labels to retain features with high predictive relevance to the classes while eliminating redundant or irrelevant features. This is based on the premise that a high-quality feature subset is one where features are both predictive of the class and non-redundant with each other Gennari et al. (1989); Roffo (a). Unsupervised learning, devoid of class labels, presents a more challenging scenario for feature selection, emphasizing the need for innovative approaches to identify relevant features without explicit class information.
@@ 2 Feature Selection Techniques
Relief-F Liu and Motoda (2008) exemplifies a supervised, iterative approach to FS, assessing feature quality by their capacity to distinguish between neighboring samples. Its effectiveness notwithstanding, Relief-F’s performance is contingent upon data availability and does not adequately address feature redundancy.
@@
SVM-RFE Guyon et al. (2002), an embedded FS method, systematically eliminates features using a backward elimination strategy, prioritizing those that significantly contribute to sample separation via a linear SVM. Despite its strengths, SVM-RFE shares Relief-F’s limitations in handling redundant features.
@@
The Fisher method Gu et al. (2012) represents a fast, filter-based FS strategy, calculating feature scores based on the ratio of interclass separation to intraclass variance. Similarly, mutual information (MI) Zaffalon and Hutter (2002) evaluates features based on the mutual information between the distribution of feature values and class membership. Both methods independently assess features before aggregating the top-ranked ones for final selection.
@@
Infinite Latent Feature Selection (ILFS) Roffo et al. (2017a); Roffo (b) introduces a generative, probabilistic approach to FS, bypassing the combinatorial challenge by considering all possible feature subsets. This method prioritizes features based on a graph weighted by PLSA-inspired mixing weights, highlighting the concept of feature relevancy.
@@
Feature selection via Eigenvector Centrality (EC-FS) Roffo and Melzi (2016c); obe (2016) employs an affinity graph where features are nodes, and node importance is determined through Eigenvector Centrality. This filter method underscores the importance of a feature as a function of its neighbors’ importance, identifying central nodes as candidate features for effective classification.
@@
Addressing the challenge of unsupervised learning scenarios, Inf-FS Roffo et al. (2015a); Roffo and Melzi (2016b) and the Laplacian Score (LS) He et al. (2005) utilize graph-based methods to assess feature importance. Inf-FS assigns importance scores by considering all possible feature subsets as paths on a graph, whereas LS evaluates features based on their ability to preserve local geometric structures through a nearest neighbor graph.
@@
Recent unsupervised wrapper methods like the Dependence Guided Unsupervised Feature Selection (DGUFS) Guo and Zhu (2018) and Feature Selection with Adaptive Structure Learning (FSASL) Du and Shen (2015) highlight the significance of data interdependencies in FS. DGUFS optimizes feature subset selection by balancing the dependence on original data and cluster labels, while FSASL integrates structure learning with FS, albeit with considerable computational demands for high-dimensional data.
@@
The Least Absolute Shrinkage and Selection Operator (LASSO) Liu and Motoda (2007) incorporates a regularization process to minimize prediction error and select features with significant coefficients, exemplifying a method where FS is a consequence of model optimization.
@@
Lastly, the feature selection via concave minimization (FSV) Bradley and Mangasarian (1998) innovatively integrates FS into the training of an SVM through linear programming, demonstrating a wrapper approach that directly ties the FS process to model training efficacy.
@@ 3 Concluding Remarks
The work presented in this paper addresses the critical area of feature selection (FS) in machine learning, offering a comprehensive review of methods and introducing the Feature Selection Library (FSLib)Roffo et al. (2020, 2019); Vinciarelli et al. (2019); Roffo and Melzi (2016d); Roffo et al. (2017b); Melzi et al. (2018); Scibelli et al. (2018); Roffo and Melzi (2016e); Bukeer et al. (2019); et al. (2016); Roffo et al. (2017b, 2020); Vinciarelli et al. (2019). This library has been designed to provide researchers and practitioners with a versatile and powerful tool for feature selection, facilitating the enhancement of machine learning models through the efficient identification of relevant features.

### 12 Mar 2024 v8

@@ Abstract
The Feature Selection Library (FSLib) introduces a comprehensive suite of feature selection (FS) algorithms for MATLAB, aimed at improving machine learning and data mining tasks. FSLib encompasses filter, embedded, and wrapper methods to cater to diverse FS requirements. Filter methods focus on the inherent characteristics of features, embedded methods incorporate FS within model training, and wrapper methods assess features through model performance metrics. By enabling effective feature selection, FSLib addresses the curse of dimensionality, reduces computational load, and enhances model generalizability. The elimination of redundant features through FSLib streamlines the training process, improving efficiency and scalability. This facilitates faster model development and boosts key performance indicators such as accuracy, precision, and recall by focusing on vital features. Moreover, FSLib contributes to data interpretability by revealing important features, aiding in pattern recognition and understanding. Overall, FSLib provides a versatile framework that not only simplifies feature selection but also significantly benefits the machine learning and data mining ecosystem by offering a wide range of algorithms, reducing dimensionality, accelerating model training, improving model outcomes, and enhancing data insights.
@@ 1 Introduction
Feature selection (FS) is a critical preprocessing step in machine learning, aimed at enhancing the performance and efficiency of classification models by reducing data dimensionality and eliminating noise. This process is instrumental in refining the dataset to include only the most informative features, thereby improving the accuracy and computational efficiency of classifiers. The essence of FS lies in its ability to filter out redundant or irrelevant information, a task that becomes increasingly challenging as data complexity grows. Traditionally, feature identification has been manually conducted by experts within various learning domains. However, this manual selection process may not always identify the most relevant or non-redundant features, highlighting the necessity for automatic feature selection methods. These automated processes are adept at discarding irrelevant, redundant, and noisy data, which is crucial for enhancing learning and classification performance. The importance of feature selection is well-acknowledged across numerous fields, including machine learning, artificial intelligence, computer vision, and data mining, finding application in information retrieval, user re-identification, recommendation systems, and visual object tracking, among others [1, 2, 3, 4, 2].
@@
Within the FS domain, methodologies are broadly classified into three primary categories: filter methods, wrapper methods, and embedded methods. Filter methods prioritize the intrinsic properties of the data, operating independently of any classifier to select features. Wrapper methods involve optimizing a predictor as part of the selection process, often yielding superior results compared to filter methods but at a higher computational cost[5, 6, 7, 8, 9, 10, 11]. Embedded methods integrate the feature selection process with the model learning mechanism, making the two inseparable. This approach is unique in its interaction between feature selection and learning, exemplifying the importance of the function class under consideration in determining feature relevance.
@@
Wrapper methods employ classifiers to assess the utility of feature subsets, embedded methods integrate selection within the classifier learning process, and filter methods analyze data’s intrinsic properties irrespective of the classifier [12]. These approaches can execute two primary operations: ranking and subset selection. Ranking evaluates the importance of individual features without considering their interactions, whereas subset selection identifies the optimal feature combination for use. Although these operations can be sequential or exclusive, subset selection is inherently supervised, contrasting with the potentially unsupervised nature of ranking [13, 14, 15, 16, 17, 18, 19].
@@
Embedded methods uniquely intertwine feature selection with the learning algorithm, making them inseparable from the learning process and thereby tailoring feature selection to the specific model in use, such as Support Vector Machines [12]. This integration of selection and learning processes within embedded methods is particularly relevant in the context of modern attention and self-attention mechanisms in transformers. These mechanisms inherently perform a form of feature selection by weighting the importance of different parts of the data (e.g., tokens in natural language processing), thus aligning with the principles of embedded feature selection methods but on a data-driven, adaptive basis.
@@
Feature selection is recognized as an NP-hard problem [12], necessitating the use of heuristic or suboptimal search strategies to manage the combinatorial complexity of selecting an optimal subset from a potentially vast feature space. Filters approach this challenge by initially ranking features individually before extracting a subset, with notable examples including Inf-FS [20], EC-FS [21], MutInf [16], and Relief-F [17]. Conversely, wrapper and embedded methods evaluate and select feature subsets in a more holistic manner, with FSV [14, 15] and SVM-RFE [13] serving as key examples. Given the NP-hard nature of feature selection, with the combinatorial explosion of feature subsets to evaluate, efficient and heuristic search strategies are employed [12]. Filters prioritize individual features through ranking before extracting a subset, exemplified by methods such as Inf-FS and MutInf, while wrappers and embedded techniques evaluate and select subsets directly, with notable examples including FSV and SVM-RFE [20, 21, 16, 17, 14, 15, 13, 22].
@@
Feature selection algorithms are also distinguished by their learning context, being either supervised or unsupervised. Supervised algorithms leverage class labels to retain features with high predictive relevance to the classes while eliminating redundant or irrelevant features. This is based on the premise that a high-quality feature subset is one where features are both predictive of the class and non-redundant with each other [23, 24]. Unsupervised learning, devoid of class labels, presents a more challenging scenario for feature selection, emphasizing the need for innovative approaches to identify relevant features without explicit class information.
@@ 2 Feature Selection Techniques
Relief-F [17] exemplifies a supervised, iterative approach to FS, assessing feature quality by their capacity to distinguish between neighboring samples. Its effectiveness notwithstanding, Relief-F’s performance is contingent upon data availability and does not adequately address feature redundancy.
@@
SVM-RFE [13], an embedded FS method, systematically eliminates features using a backward elimination strategy, prioritizing those that significantly contribute to sample separation via a linear SVM. Despite its strengths, SVM-RFE shares Relief-F’s limitations in handling redundant features.
@@
The Fisher method [19] represents a fast, filter-based FS strategy, calculating feature scores based on the ratio of interclass separation to intraclass variance. Similarly, mutual information (MI) [16] evaluates features based on the mutual information between the distribution of feature values and class membership. Both methods independently assess features before aggregating the top-ranked ones for final selection.
@@
Infinite Latent Feature Selection (ILFS) [25, 26] introduces a generative, probabilistic approach to FS, bypassing the combinatorial challenge by considering all possible feature subsets. This method prioritizes features based on a graph weighted by PLSA-inspired mixing weights, highlighting the concept of feature relevancy.
@@
Feature selection via Eigenvector Centrality (EC-FS) [21, 27] employs an affinity graph where features are nodes, and node importance is determined through Eigenvector Centrality. This filter method underscores the importance of a feature as a function of its neighbors’ importance, identifying central nodes as candidate features for effective classification.
@@
Addressing the challenge of unsupervised learning scenarios, Inf-FS [20, 3] and the Laplacian Score (LS) [28] utilize graph-based methods to assess feature importance. Inf-FS assigns importance scores by considering all possible feature subsets as paths on a graph, whereas LS evaluates features based on their ability to preserve local geometric structures through a nearest neighbor graph.
@@
Recent unsupervised wrapper methods like the Dependence Guided Unsupervised Feature Selection (DGUFS) [29] and Feature Selection with Adaptive Structure Learning (FSASL) [30] highlight the significance of data interdependencies in FS. DGUFS optimizes feature subset selection by balancing the dependence on original data and cluster labels, while FSASL integrates structure learning with FS, albeit with considerable computational demands for high-dimensional data.
@@
The Least Absolute Shrinkage and Selection Operator (LASSO) [31] incorporates a regularization process to minimize prediction error and select features with significant coefficients, exemplifying a method where FS is a consequence of model optimization.
@@
Lastly, the feature selection via concave minimization (FSV) [14] innovatively integrates FS into the training of an SVM through linear programming, demonstrating a wrapper approach that directly ties the FS process to model training efficacy.
@@ 3 Concluding Remarks
The work presented in this paper addresses the critical area of feature selection (FS) in machine learning, offering a comprehensive review of methods and introducing the Feature Selection Library (FSLib)[39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 43, 39, 41]. This library has been designed to provide researchers and practitioners with a versatile and powerful tool for feature selection, facilitating the enhancement of machine learning models through the efficient identification of relevant features.